OpenStudy (chrisdbest):

Rewrite with only sin x and cos x. cos 3x (A) cos x - 4 cos x sin2x (B)-sin3x + 2 sin x cos x (C)-sin2x + 2 sin x cos x (D)2 sin2x cos x - 2 sin x cos x

OpenStudy (welshfella):

Hint cos 3x = cos(2x + x)

OpenStudy (chrisdbest):

Ye s I know that.

OpenStudy (chrisdbest):

I have a headache from a slip and fall accident and just don't feel like working. I will medal you if you have the correct answer.

OpenStudy (welshfella):

but I dont think we can use that anyway because cos ( a + b) = cosa cosb - sina sinb

OpenStudy (chrisdbest):

-_-

OpenStudy (chrisdbest):

So can someone please tell me the answer

OpenStudy (welshfella):

Its not C because that simplifies to -sin 2x + sin2x = 0

OpenStudy (chrisdbest):

ohkay

OpenStudy (chrisdbest):

I know it's not D because I chose that and it was wrong

OpenStudy (welshfella):

I cant convert cos 3x into any of the options . I think there might be an error in the question. I'll plug in x = 20 degrees and see if any of them work out to 0.5 (because cos 60 = 0.5)

OpenStudy (chrisdbest):

try B then

OpenStudy (welshfella):

- sin 60 + 2 cos 20 sin 20 = -0.2232 not that one

OpenStudy (welshfella):

nor A - none of them equal cos 3x

OpenStudy (welshfella):

thers a problem with the question

OpenStudy (chrisdbest):

That's the beauty of Virtual School

OpenStudy (chrisdbest):

I know it's not D or B, because I just got it wrong and can't do it over again. I have been told it's not A or C by an elite, and I KEEP GETTING THE QUESTION WRONG!!!!